A 2.8 kg mass is lifted from a height of 2.5 m to a height of 8.0 m in 3.3 s. the power of the lift is
1
Expert's answer
2020-06-15T10:35:51-0400
"P=\frac{mg(h_2-h_1)}{t}"

"P=\frac{(2.8)(9.8)(8-2.5)}{3.3}=46\ W"